You have 1.00g of your organic product dissolved iOrganic Chemistry
Answer #1

Case Study:-(1) Weight of Organic Product =1 g (2)Volume of Solvent (water)=50 ml (3)Distribution coefficient of this compound in Ether and water =5

Answer:-

Distribution Coefficient is the ratio of Concentration of a compound in the mixture of two immissible phase at equilibrium.this also measure the difference in solubility of the compound in these two phase.In this case we have two immissible phase one is organic phase and another is aqueous phase .

Here we are define organic phase is mixture of organic compound with water and Aqueous phase is pure ether.

(1)Organic phase :-1 g of organic compound dissolve in 50 ml of water solution thus volume of organic compound =50 ml

(2)Aqueous Phase(Ether) :-(a)In first case volume of ether =60 ml(b)In second case volume of ether =20ml

S aq\leftrightharpoons Sorg

So Distribution Coefficient (D)=(Concentraion of solid in organic phase )/(Concentraion of solid in aqueous phase)

D=5

Now Concetration =moles of compound /Volume of solution

We derive the equation

Fraction of organic compound(q org) =DVaq / (D V org+ V aq)

(1) Fraction of organic compound(q org)1=(5x60)/((5x50)+(60))=300/310=0.96

So in first answer organic compound is extract 96 %

(2)Fraction of organic compound(q org)2=(5x20)/((5x50)+(20))=100/270=0.3703

So in second answer organic compound is extract 37.03 % on each 20 ml of ether